OpenStudy (anonymous):

Two reactions have the same rate at 30 degree C. The first reaction doubles its rate when heated to 40 degree C. The activation energy of the second reaction is four times that of the first. Compare the rates of these two reactions at 100 degree C

OpenStudy (aaronq):

use the two-point form of the Arrhenius equation: \(ln\dfrac{k_1}{k_2}=\dfrac{E_a}{R}(\dfrac{1}{T_1}-\dfrac{1}{T_2})\)

OpenStudy (anonymous):

thank you :)